# Prosody Analysis of Audiobooks

### v2

@@ Abstract
Recent advances in text-to-speech have made it possible to generate natural-sounding audio from text. However, audiobook narrations involve dramatic vocalizations and intonations by the reader, with greater reliance on emotions, dialogues, and descriptions in the narrative. Using our dataset of 93 aligned book-audiobook pairs, we present improved models for prosody prediction properties (pitch, volume, and rate of speech) from narrative text using language modeling. Our predicted prosody attributes correlate much better with human audiobook readings than results from a state-of-the-art commercial TTS system: our predicted pitch shows a higher correlation with human reading for 22 out of the 24 books, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books. Finally, we present a human evaluation study to quantify the extent that people prefer prosody-enhanced audiobook readings over commercial text-to-speech systems.
1 Introduction
@@
Text-to-speech (TTS) technology has made it possible to create machine-generated readings of text. With recent advances in the use of deep learning for TTS models (Oord et al., 2016; Arık et al., 2017), the ability to produce natural-sounding audio closer to human speech has also improved. But there remains a sizable performance gap, especially in the domains like story narration which depend on a global interpretation of a book-length text. Audio books are often recorded by professional actors, involving dramatic vocalizations and intonations by the reader to emphasize emotions, dialogues, and environment and character descriptions in the narrative. Although machine-generated audio sounds human-like in nature, it does not account for these elements in narrative readings, which show much variation in intonation and prosody.
In this paper, we analyze audio books from an NLP-perspective, with two distinct objectives in mind. First, we study how higher-order NLP analysis of novels (e.g. character identification, quote/dialog analysis, and narrative flow analysis) might be used to build better TTS systems.
Second, we are interested as audio books as a source of annotations for narrative texts. One of the challenges of NLP on book-length documents is the cost of annotation: reading a novel is a 10-15 hour commitment, making it cost prohibitive to do human annotation for special-purpose tasks on a large corpora of books. We see human-recorded audio books as a potential solution here: each human recording implicitly contains information about the contents of the texts, which analysis of the audio may enable us to programmatically extract.
State-of-the-art TTS systems do not approach the dynamic range and expressiveness of human readers. Machine-generated audio consistently stays within a small range of pitch and volume, indicative of a flatter reading. In this work, we address the task of generating higher-quality audiobooks closer to human readings in nature, by predicting prosody attributes from the text. Our main contributions in this work include:
  • Character-level analysis of audiobook reader behaviors — We quantify the extent to which the gender properties of characters are reflected in audiobook readings. Specifically, in 21 of 31 books where the two lead characters differ in gender, readers used lower pitch and higher volume to portray the male character. In 52 of 62 books with significant dialog for both genders, reader pitch was higher for female characters. Further, readers generally use lower pitch in narrative regions rather than dialog, independent of gender.
  • Models for audiobook prosody prediction — We address the task of predicting prosody attributes given a narrative text, training a variety of models on audiobook readings to predict prosody attributes (pitch, volume, and rate) at the sentence-level and phrase-level. Best results for all three prosody attributes came using a LSTM architecture, built upon MPNet embeddings, operating at the phrase level. Our predicted prosody attributes correlate much better with human audiobook readings than results from a state-of-the-art commercial TTS system (Google Cloud Text-to-Speech). The dot plot of Figure 1 presents the human-TTS correlation for two systems, on a test set of the first chapters of 24 books. Our predicted pitch attribute shows a higher correlation with human reading for 22 out of the 24 books, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books.
@@
  • Human evaluation study for text-to-speech audiobook reading — The prosody attributes predicted by our models from text can be directly incorporated to enhance TTS systems through the Speech Synthesis Markup Language (SSML). We conduct a human evaluation study to understand whether humans prefer to hear our more expressive SSML-enhanced text-to-speech audio over the commercial Google Cloud Text-to-Speech, a strong baseline. Results here are inconclusive, but a small majority (12 of 22) subjects preferred our SSML-enhanced readings.
@@
Over the course of this study we have developed a dataset of phrase-aligned books and audiobooks, comprising 1806 chapters across 93 novels, with their corresponding human-read audiobooks and phoneme-level alignment between the text and audio. This dataset will be made publicly available at www.anonymous.com upon acceptance.
@@
This paper is organized as follows. Section 2 presents related work associated with text-to-speech synthesis, including an overview of the Speech Synthesis Markup Language (SSML) we use in our experiments. Section 3 describes our methods for the critical phase of alignment between the textual representation of novels and associated audio book recordings. Section 4 presents our analysis of reader behavior, quantifying the degree to which voices change with a character’s gender. We develop prosody prediction models from text in Section 5, culminating in a human preference study reported in Section 6. Future directions for our work are discussed in Section 7.
2 Related Work
NLP for books:
Recent work focused on performing NLP tasks on novels (Bamman et al., 2014; Brooke et al., 2015; StonyBook, n.d.) has facilitated the analysis and downstream use of book artifacts (Pethe et al., 2020; Kim et al., 2020; Wilmot and Keller, 2020, 2021; Papalampidi et al., 2020; Otake et al., 2020). These systems employ CoreNLP (Manning et al., 2014), Stanza (Qi et al., 2020), and SpaCy (Honnibal and Montani, 2017) for NLP annotation tasks including tokenization, lemmatization, part-of-speech tagging, dependency tagging, named entity recognition, coreference resolution, and quote extraction and attribution.
Approaches to character identification in novels involves clustering variants of character names (Davis et al., 2003; Elson et al., 2010; Vala et al., 2015; Jahan and Finlayson, 2019). (Groza and Corde, 2015) use a folktale-specific ontology based approach. The CHARLES tool (Vala et al., 2016) provides a collaborative interface for the manual annotation of characters. (Dekker et al., 2019) compare four different approaches for character identification by evaluating against a manually annotated dataset. Further, methods to generate character embeddings and analyze relationships between characters (Inoue et al., 2022; Iyyer et al., 2016; Xanthos et al., 2016; Azab et al., 2019; Brahman et al., 2021) provide additional information about characters apart from their mentions in the book.
Genre identification (Worsham and Kalita, 2018), tracking emotion in narratives (Mohammad, 2011), and quote identification and attribution (O’Keefe et al., 2012; Zhang et al., 2003) are additional tasks that provide context about events occurring in the book narrative.
These artifacts obtained from analysis of books contain information that influences the way the readings of these books can be customized as per the type of content, mood, characters’ nature, and context of dialogue.
Text to Speech:
Speech synthesis or text-to-speech (TTS) is the technology to produce artificial human speech. With recent advances in deep learning, neural network-based TTS models have become the dominant approach. WaveNet (Oord et al., 2016) is a neural vocoder that directly generates time domain waveforms from linguistic features without the use of acoustic features, and it outperforms traditional statistical parametric synthesizers. Tacotron(Wang et al., 2017) uses a single neural network to directly generate spectrograms from a sequence of characters without the use of linguistic features. (Shen et al., 2018) extend this work by combining it with a modified WaveNet to improve voice quality. Naturalness and robustness of speech synthesis are improved based on this work by transferring prosody and speaking style from reference signals (Skerry-Ryan et al., 2018), using reference encoder and style tokens (Wang et al., 2018), hierarchical latent variables (Hsu et al., 2019), multi-reference encoder (Bian et al., 2019), and duration predictors (Yu et al., 2019; Shen et al., 2020). (Weiss et al., 2021) propose Wave-tacotron, an end-to-end model that generates waveform directly from text. Tacotron model is further improved by changing sequence generation from autoregressive to non-autoregressive (Elias et al., 2021b, a).
DeepVoice (Arık et al., 2017) is a CNN-based model that uses independent neural networks in place of traditional TTS components. (Arik et al., 2017) extend this work by adding multi-speaker support using Tacotron-style model. DeepVoice 3 proposed by (Ping et al., 2017) has a fully-convolutional attention-based encoder-decoder architecture that supports larger datasets and faster training. FastSpeech (Ren et al., 2019) is a non-autoregressive Transformer-based parallel TTS. It uses knowledge distillation and is able to fully utilize GPU and synthesize spectrograms faster than autoregressive models. Its monotonic constraint on alignment also reduces synthesis failure. FastSpeech 2 is proposed by (Ren et al., 2020), which solves the one-to-many mapping problem better by conditioning on supervised fundamental frequency and energy. FastPitch (Łańcucki, 2021) uses a similar approach. Lightweight TTS models (Luo et al., 2021; Nguyen et al., 2021; Vainer and Dušek, 2020) are efficient to deploy on mobile and edge devices.
More recent works on TTS include Pitchtron (Jung and Kim, 2020), which proposes two prosody transfer models to generate speech with rich prosody. (Sini, 2020) investigates different characteristics that contribute to the expressiveness of a voice. (Sivaprasad et al., 2021) present a TTS system that allows user to generate speech with chosen emotion from a continuous and interpretable space of arousal and valence.
Some automated audiobook generation vendors include DeepZen, Speechki, Speechkit, Descript, Pozotron, Scribe Audio, Fabula, Respeecher, and Accessibility.
2.1 SSML Overview
@@
Speech Synthesis Markup Language (SSML) (Taylor and Isard, 1997) is a markup language that provides a standard way to mark up text for the generation of synthetic speech, allowing for more fine-grained control over the generated audio. In this work, we are specifically interested in the <prosody> tag, which can be used to control three different attributes: pitch, volume, and rate of speech:
@@
  • Pitch — The pitch can be increased or decreased by N percent relative to the default value, by setting the pitch attribute as "+N%" or "-N%" respectively.
  • Volume — The volume can be increased or decreased by N decibels relative to the default value, by setting the volume attribute as "+NdB" or "-NdB" respectively.
  • Rate — The rate of speech can be specified as N percent of the default value by setting the rate attribute to "N%".
3 Audio/Text Alignment Methods
With the end goal of enhancing prosody in automatically generated audiobooks, we present a dataset of pairs of books and their corresponding audiobooks from the Project Gutenberg dataset, along with the extracted prosody data from the human-read audiobooks.
The dataset contains a total of 1806 chapters across 93 books, with a mean of 19.42 chapters per book. Generally each Gutenberg audiobook is read by a single narrator, but there are exceptions, so we performed our analysis on a per-chapter basis. We used a mix of narrator annotation and audio speaker identification techniques to establish narrator consistency.
We perform the following data processing steps to generate aligned data and extract the audio attributes.
  • Transcription — We first generate a text transcript of each audiobook. We split the audiobook into fragments using silence for segmentation. For each segment, we attempt to generate a transcript using Google Speech Recognition  [^2] https://pypi.org/project/SpeechRecognition/. Note that this transcript is noisy, and cannot be used directly for further tasks. We use this noisy transcript to align with the appropriate chapter, and use this text for downstream tasks.
@@ 3 Audio/Text Alignment Methods
  • Audio Attribute Extraction — To extract the pitch (fundamental frequency in Hz) and the volume (intensity in dB), we use the Python library Parselmouth [^4] https://parselmouth.readthedocs.io/en/stable/, with the timestep set to 0.01 second (compatible with the granularity of the alignment output from Gentle).
    To extract the rate of speech, we perform the following steps: For each phoneme, we first compute the z-scores of the durations of all occurrences. Then for each audio segment (phrase / sentence), we compute the mean of the phoneme duration z-scores, and compute the z-score of the resultant means.
  • Sentence Segmentation — For each sentence, we generate a constituency parse tree using the Berkeley Neural Parser (Kitaev and Klein, 2018) with SpaCy (Honnibal and Montani, 2017). We then split the sentence into phrases, using the constituent sentence (S) components and punctuation.
4 Audiobook Reader Behavioral Analysis
@@
We also compare pitch and volume for the most frequent and the second most frequent characters in each book. Figure 3(b) shows pitch and volume of character pairs colored by gender. When the two main characters are of opposite gender, female character has a higher pitch and a lower volume compared to male character as most red dots are below the regression line, and most purple dots are above on both plots. As shown in Table 1, when the two main characters are of opposite gender, we have a 67.74% chance of distinguishing female from male based on pitch or volume.
Prosody
Gender of 1st and
2nd character
1st > 2nd 1st < 2nd
Pitch Female, male 13 5
Male, female 5 8
Male, male 19 18
Female, female 8 5
Volume Female, male 6 12
Male, female 9 4
Male, male 22 15
Female, female 4 9
Rate Female, male 8 10
Male, female 5 8
Male, male 17 20
Female, female 9 4
Table 1: Number of character pairs above and below the regression line for pitch, volume, and rate of speech.
Finally, we compare the pitch and volume for dialogue utterances averaged separately across all the male and female characters in a book. For this comparison, we restrict our analysis to 62 books in which male and female characters each have at least 100 combined word utterances.
Figure 4: Mean pitch for dialog of male and female characters. Female characters are voiced with higher pitch.
For 52 out of 62 books, the mean pitch is higher for female character dialogue as compared to male (binomial p=3e-8). For 32 out of 62 books, the mean intensity (volume) is higher for male character dialogue as compared to female (binomial p=0.449).
5 Predicting Prosody Attributes
@@
5.1 Text Input Representations
We experiment with three input embeddings:
  • TF-IDF representations: We use TF-IDF vectorization, limiting the vocabulary to 1,000 words to generate 1000-dimensional representations of the input text.
  • Mean-pooled GloVe embeddings: We compute the 300-dim average word embedding of the input text using the CommonCrawl pre-trained GloVe model (840B tokens).
  • MPNet embeddings: We compute 768-dimensional embeddings of the input text using MPNet (Song et al., 2020) (all-mpnet-base-v2).
We use Sentence-Transformers  [^5] https://www.sbert.net/ to compute the GloVe and MPNet embeddings. We process the audio attributes extracted as described in Section 3 for this task. For each text segment in the input, we consider the values for the three prosody attributes (pitch, volume, rate) in the corresponding audio segment. We compute a single value for each attribute per text segment, and compute the z-scores across all segments in the chapter.
5.2 Models
@@
5.3 Sentence-level Attribute Prediction
We perform the prediction task at the sentence-level granularity. In this section, we present the test metrics for baseline models, where we use only a single sentence representation as the input, and predict a single prosody attribute (pitch/volume/rate) as the output.
Embedding Model Pitch Volume Rate
TF-IDF LinReg 0.9217 0.9054 0.8439
MLP (5, 5) 0.9165 1.0000 1.0000
MLP (10, 10) 0.9636 0.9342 0.8034
MLP (20, 20) 1.0554 1.0127 0.8605
GloVe LinReg 0.9103 0.8900 0.8578
MLP (5, 5) 0.9088 0.8955 0.8212
MLP (10, 10) 0.9208 0.8967 0.8396
MLP (20, 20) 0.9749 0.9395 0.8734
MPNet LinReg 0.8667 0.8440 0.7811
MLP (5, 5) 0.8785 0.8457 0.7996
MLP (10, 10) 0.9403 0.8888 0.8519
MLP (20, 20) 1.0642 0.9742 0.8899
Table 2: Test MSE for sentence-level prosody attributes
Table 2 shows the test metrics for the respective baseline models. As the sentence embedding quality improves, we expect the error to reduce. The test metrics are in accordance with this, as MPNet embeddings have the least prediction error across all three attributes.
5.4 Phrase-level Attribute Prediction
A drawback of the sentence-level model is the inconsistency in sentence length. Sentences may range from 1 to more than 500 tokens. In order to introduce consistency, we use phrase representations as input. We segment sentences into phrases using the method described in Section 3.
Table 3 shows the test metrics for the phrase-level baseline models. These models also show similar trends in the results, with MPNet embeddings showing the best performance.
As the MSE for the phrase-level models is lower than that of the sentence-level models, we present results from the phrase-level models subsequently.
Embedding Model Pitch Volume Rate
TF-IDF LinReg 0.9217 0.8789 0.8439
MLP (5, 5) 0.9086 0.8594 0.8145
MLP (10, 10) 0.9104 0.8623 0.8058
MLP (20, 20) 0.9155 0.8700 0.8073
GloVe LinReg 0.9197 0.8644 0.9225
MLP (5, 5) 0.9095 0.8462 0.8717
MLP (10, 10) 0.9037 0.8449 0.8597
MLP (20, 20) 0.9123 0.8496 0.8569
MPNet LinReg 0.8803 0.8048 0.8123
MLP (5, 5) 0.8661 0.7812 0.7763
MLP (10, 10) 0.8675 0.7814 0.7791
MLP (20, 20) 0.8955 0.7933 0.7987
Table 3: Test MSE for phrase-level prosody attributes
5.5 Joint Prediction
@@
Using only a single phrase embedding as input (with no contextual information about the neighboring phrases), we predict the three attributes. Table 4 shows the test MSE for the three attributes. These models show a slight improvement over the best model results from Section 5.4.
@@
Table 4: Test MSE for non-contextual joint prediction of prosody attributes
5.5.1 Contextual Prediction
We augment the feature space with embeddings of context phrases before and after each phrase, and jointly predict the three attributes.
Embedding Model Pitch Volume Rate
GloVe LinReg 0.9121 0.8513 0.9154
MLP (5, 5) 0.9081 0.8422 0.8836
MLP (10, 10) 0.9056 0.8387 0.8645
MLP (20, 20) 0.9070 0.8439 0.8856
MPNet LinReg 0.8620 0.7779 0.7813
MLP (5, 5) 0.8659 0.7772 0.7852
MLP (10, 10) 0.8594 0.7671 0.7751
MLP (20, 20) 0.9381 0.7831 0.8586
Table 5: Test MSE for contextual joint prediction of prosody attributes
Table 5 shows the test MSE for the three attributes. This is again a slight improvement over non-contextual prediction (Table 4).
5.5.2 Sequential Prediction
We use a neural network (LSTM architecture and training described in Section 5.2) to predict the prosody attributes for all input. We experiment with sequence lengths 2 and 3.

### v3

@@ Abstract
Recent advances in text-to-speech have made it possible to generate natural-sounding audio from text. However, audiobook narrations involve dramatic vocalizations and intonations by the reader, with greater reliance on emotions, dialogues, and descriptions in the narrative. Using our dataset of 93 aligned book-audiobook pairs, we present improved models to predict prosody (pitch, volume, and rate of speech) from narrative text using language modeling. Our predicted prosody attributes correlate much better with human audiobook readings than results from a state-of-the-art commercial TTS system: our predicted pitch shows a higher correlation with human reading for 22 out of 24 books in the test set, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books. Finally, we present a human evaluation study to quantify the extent that people prefer prosody-enhanced audiobook readings over default commercial text-to-speech systems.
Index Terms:
prosody attribute prediction, text to speech, character embedding
I Introduction
@@
In this paper, we analyze audio books from an NLP perspective, with two distinct objectives in mind. First, we study how higher-order NLP analysis of novels (e.g. character identification, quote/dialog analysis, and narrative flow analysis) might be used to build better TTS systems. Second, we are interested as audio books as a source of annotations for narrative texts. One of the challenges of NLP on book-length documents is the cost of annotation: reading a novel is a 10-15 hour commitment, making it cost-prohibitive to do human annotation for special-purpose tasks on a large corpora of books. We see human-recorded audio books as a potential solution here: each human recording implicitly contains information about the contents of the texts, which we can programmatically extract by analysing the audio. Our main contributions in this work include:
  • Character-level analysis of audiobook reader behaviors — We quantify the extent to which the gender properties of characters are reflected in audiobook readings. Specifically, in 21 of 31 books where the two lead characters differ in gender, readers used lower pitch and higher volume to portray the male character. Further, readers generally use lower pitch in narrative regions rather than dialog, independent of gender.
  • Models for audiobook prosody prediction — We address the task of predicting prosody attributes given a narrative text, training a variety of models on audiobook readings to predict prosody attributes (pitch, volume, and rate). The dot plot of Figure 1 presents the human-TTS correlation for our system using a LSTM architecture built upon MPNet embeddings, and a state-of-the-art commercial TTS system (Google Cloud Text-to-Speech) on a test set of the first chapters of 24 books. Our predicted pitch attribute shows a higher correlation with human reading for 22 out of the 24 books, while our predicted volume attribute proves more similar to human reading for 23 out of the 24 books.
@@
  • Human evaluation study for text-to-speech audiobook reading — We conduct a human evaluation study to understand whether humans prefer to hear our more expressive text-to-speech audio enhanced with Speech Synthesis Markup Language (SSML) over the default commercial Google Cloud Text-to-Speech, a strong baseline. Results are inconclusive, but a small majority (12 of 22) subjects preferred our SSML-enhanced readings.
@@
Over the course of this study, we have developed a dataset of sentence-aligned books and audiobooks comprising 1806 chapters across 93 novels, with their corresponding human-read audiobooks and sentence-level alignment between the text and audio. This dataset has been made publicly available at https://github.com/sbu-dsl/prosody-analysis-of-audiobooks.
@@
This paper is organized as follows. Section II presents related work associated with NLP for books and an overview of the Speech Synthesis Markup Language (SSML) we use in our experiments. Section III describes our methods for the critical phase of alignment between the textual representation of novels and associated audio book recordings. Section IV presents our analysis of reader behavior, quantifying the degree to which voices change with a character’s gender. We develop prosody prediction models from text in Section V, culminating in a human preference study reported in Section VI. Future directions for our work are discussed in Section VII.
II Related Work
II-A NLP for books:
Recent work focused on performing NLP tasks on novels [1, 2, 3] has facilitated the analysis and downstream use of book artifacts [4, 5, 6, 7, 8]. These systems employ CoreNLP [9], Stanza [10], and SpaCy [11] for NLP annotation tasks. Further, methods to generate character embeddings and analyze relationships between characters [12, 13, 14, 15] provide additional information about characters apart from their mentions in the book.
II-B SSML Overview
@@
Speech Synthesis Markup Language (SSML) [17] is a markup language that provides a standard way to mark up text for the generation of synthetic speech, allowing for more fine-grained control over the generated audio. In this work, we are specifically interested in the <prosody> tag, which can be used to control three different attributes: pitch, volume, and rate of speech. We explore the role of audio in providing additional context and emotional cues [18]. An example of SSML is shown in Figure 2.
@@
III Audio/Text Alignment Methods
We present a dataset of pairs of books and their corresponding audiobooks from the Project Gutenberg dataset, along with the extracted prosody data from the human-read audiobooks. The dataset contains a total of 1806 chapters across 93 books. We perform the following data processing steps to generate aligned data and extract the audio attributes.
  • Transcription — We first split the audiobook into fragments using silence for segmentation. For each segment, we attempt to generate a transcript using Google Speech Recognition  [^2] https://pypi.org/project/SpeechRecognition/. Note that this transcript is noisy, and we only use it to align with the appropriate chapter.
@@ III Audio/Text Alignment Methods
  • Audio Attribute Extraction — To extract the pitch (fundamental frequency in Hz) and the volume (intensity in dB), we use the Python library Parselmouth [^4] https://parselmouth.readthedocs.io/en/stable/, with the timestep set to 0.01 second (compatible with the granularity of the alignment output from Gentle). To extract the rate of speech, we first compute the z-scores of the durations of all occurrences for each phoneme. Then we compute the mean of the phoneme duration z-scores for each audio segment (phrase / sentence), and the z-score of the resultant means.
  • Sentence Segmentation — For each sentence, we generate a constituency parse tree using the Berkeley Neural Parser [17] with SpaCy [11]. We then split the sentence into phrases, using the constituent sentence (S) components and punctuation.
IV Audiobook Reader Behavioral Analysis
@@
We compare pitch and volume for the most frequent and the second most frequent characters in each book. Figure 3(b) shows pitch and volume of character pairs colored by gender. When the two main characters are of opposite gender, female character has a higher pitch and a lower volume compared to male character as most red dots are below the regression line, and most purple dots are above on both plots.
We also compare the pitch and volume for dialogue utterances averaged separately across all the male and female characters in a book. For this comparison, we restrict our analysis to 62 books in which male and female characters each have at least 100 combined word utterances. For 52 out of 62 books, the mean pitch is higher for female character dialogue as compared to male (binomial p===3e-8). For 32 out of 62 books, the mean intensity (volume) is higher for male character dialogue as compared to female (binomial p===0.449).
V Predicting Prosody Attributes
@@
V-A Text Input Representations
We experiment with three input embeddings: TF-IDF representations, Mean-pooled GloVe embeddings computed from the CommonCrawl pre-trained GloVe model (840B tokens), and MPNet embeddings [19] (all-mpnet-base-v2). We compute a single value for each prosody attribute per text segment, and compute the z-scores across all segments in the chapter.
@@ V-A Text Input Representations
V-B Joint Prediction
@@
Using only a single phrase embedding as input (with no contextual information about the neighboring phrases), we predict the three attributes. Table I shows the test MSE for the three attributes.
  • LinReg: We use linear regression as the baseline model, with the default parameters specified in sklearn [^5] https://scikit-learn.org/stable/modules/generated/sklearn.linear_model.LinearRegression.html.
  • MLP: We use the MultiLinear Perceptron with the default parameters specified in sklearn [^6] https://scikit-learn.org/stable/modules/generated/sklearn.neural_network.MLPRegressor.html, with hidden layer sizes of (5,5), (10,10) and (20,20) respectively.
@@
TABLE I: Test MSE for non-contextual joint prediction of prosody attributes
V-C Sequential Prediction
We use an LSTM model with input sequence length 2 or 3, with the following architecture: a Bidirectional LSTM layer (size 40), tanh-activation Dense layer (size 20), linear-activation (size 3). We use Mean Squared Error (MSE) as the loss function. We use a validation split of 15%, train for 30 epochs, and select the model with the least validation loss.